OpenStudy (anonymous):

Find the common difference of the sequence shown. 1/2, 1/4, 0, ... -1/8 -1/4 -1/2

OpenStudy (anonymous):

-1/4 right?

Parth (parthkohli):

-1/4, notice the pattern, you'll get it. 1/2 - 1/4 = 1/4 1/4 - 1/4 = 0

OpenStudy (savvy):

that is just the 2nd term - 1st term...which is 1/4

OpenStudy (anonymous):

Yes, (-1/4) is correct. By subtracting (or adding) (-1/4) to each term, you find the next term of the sequence.

OpenStudy (anonymous):

thanks guys and girls

OpenStudy (anonymous):

You're quite welcome! :)